OpenStudy (anonymous):

Which of the following represents the graph of the equation y = 5x – 4?

OpenStudy (anonymous):

which way would the line go?

OpenStudy (anonymous):

Okay, let's convert it into a function.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x) = 5x - 4 We can do this because y = f(x).

OpenStudy (anonymous):

Now you need to choose two values for x. What would you like to use?

OpenStudy (anonymous):

i dont know wat u mean?

OpenStudy (anonymous):

Pick two numbers between 1 and 10.

OpenStudy (anonymous):

3 and 8

OpenStudy (anonymous):

It has a slope of 5/1 and at the b intersect, which is -4 go to the b intersect and then rise 5 and run 1. That will be your line

OpenStudy (anonymous):

Okay, cool. f(x) = 5x - 4 Let x = 3 f(3) = 5(3) - 4 Let x = 8 f(8) = 5(8) - 4

OpenStudy (anonymous):

MDoodler, please read the Code of Conduct and note that you are not supposed to give answers, but help people find them.

OpenStudy (anonymous):

Dustin, would you like me to explain how to do it, or would you rather use his answer?

OpenStudy (anonymous):

that is not my answer anyways

OpenStudy (anonymous):

I figured.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

jcpd I am not giving answer but explaining to the person one way to solve the problem

OpenStudy (anonymous):

Okay, back to what I was saying.

OpenStudy (anonymous):

f(x) = 5x - 4 Let x = 3 f(3) = 5(3) - 4 Let x = 8 f(8) = 5(8) - 4

OpenStudy (anonymous):

So now you have to solve the equations. Do you know how to do that?

OpenStudy (anonymous):

@Dustin15

OpenStudy (anonymous):

how do you know what y is?

OpenStudy (anonymous):

We're solving for y. :)

OpenStudy (anonymous):

Remember y = f(x)

OpenStudy (anonymous):

you do know i am graphing right?

OpenStudy (anonymous):

So when we solve for f(3) and f(8) we are solving for y when x equals 3 and when x equals 8

OpenStudy (anonymous):

You are graphing a line.

OpenStudy (anonymous):

Trust me. :D

OpenStudy (anonymous):

f(3) = 5(3) - 4 f(3) = 15 - 4 {multiply 5 by 3} f(3) = 11 {subtract 4 from 15} y = 11 So when x = 3, y = 11. Therefor, an ordered pair on this line is: (3,11)

OpenStudy (anonymous):

Want me to help you find the next ordered pair?

OpenStudy (anonymous):

He explains it well.